# Learning Contrastive Feature Representations for Facial Action Unit Detection

### v1

@@ III Method
III-A Contrastive Learning
To tackle the first challenge, we propose a contrastive learning framework that encourages the encoder to focus on capturing differences of each AU type instead of the pixel-level information of the whole face. By leveraging the relationships between samples and their corresponding AUs, the framework promotes the extraction of meaningful and discriminative features for AU detection. We investigate the feature representation of a specific AU, denoted as $x$. For AU detection which belongs to multi-label binary classification problem, the cross-entropy loss emerges as the primary choice for training models. This loss captures the discrepancy between predicted and ground truth labels, providing a principled approach to guide the learning process. Specifically, for a specific AU, the classification loss is calculated by comparing the features of samples with their corresponding labels. This enables the model to optimize its parameters with respect to the task of accurately classifying AUs. Mathematically, the typical cross-entropy loss can be expressed as follows:
$\mathcal{L}_{\textsc{ce}}=y_{i}\log\hat{y}_{i}+(1-y_{i})\log(1-\hat{y}_{i})$ (1)
where $y_{i}$ denotes the ground truth value of occurrence (i.e., 0 or 1) of the $i$-th AU, and $\hat{y}_{i}$ is the corresponding predicted occurrence probability. This loss effectively guides the training process, encouraging the model to allocate higher probabilities to the correct class labels and lower probabilities to incorrect ones. In the case of sample $x_{i}$, the loss aims to minimize the discrepancy between predicted probabilities and the ground truth labels of the models. When the ground truth label is 1, the model is expected to assign a high probability to this class. Similarly, when the ground truth label is 0, the model should assign a low probability to this class.
However, there is variation in the occurrence rates of different AUs, with some having low rates while others have high rates. Therefore, for AU detection, we usually use the weighted cross-entropy loss [11] ${L}_{wce}$, which is formulated as:
${L}_{wce}=-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}[y_{i}log\hat{y}_{i}+(1-{y}_{i})log(1-\hat{y}_{i})]$ (2)
where $w_{i}$=$\frac{(1/r_{i})n_{au}}{\sum_{i=1}^{n_{au}}(1/r_{i})}$ is the weight of the $i$-th AU determined by its occurrence frequency for balanced training. It permits reducing the impact of loss values associated with those AUs that are more commonly activated in the training set. Consequently, loss values arising from less frequently inactivating AUs are given greater significance during the training process. $r_{i}$ represents the occurrence frequency of the $i$-th AU, and $n_{au}$ is the total number of AUs.
The classification methods based on cross-entropy loss are generally applicable to most classification tasks. However, they suffer from a common issue of being overly constrained, which can lead to overfitting and limited generalization. This constraint becomes particularly problematic in the context of AU detection, where different AUs share a common encoder. For each AU of every sample, the predicted output $\hat{y}$ must satisfy the aforementioned condition, placing a significant burden on the fitting capability of the encoder. To address this challenge, we introduce a contrastive learning framework in the task of AU detection:
$\mathcal{L}_{wnce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{+}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (3)
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))}]\}$
In the equation, $f(\cdot)$ serves as an encoder, mapping the features to a $d$-dimensional vector. The positive sample of $x$, denoted as $x^{+}$, corresponds to samples belonging to the same class as $x$. Typically, $x^{+}$ is obtained by applying various semantic-invariant operations such as random cropping or Gaussian blurring to $x$. On the other hand, $x^{-}$ represents negative samples of $x$, which are samples from different classes compared to $x$. $\text{sim}(f(x),f(x^{+}))$ measures the similarity between $x$ and its positive example $x^{+}$, while $\text{sim}(f(x),f(x^{-}_{j}))$ measures the similarity between the sample and its negative example. The term ${p}_{d}$ represents the data distribution, ${p}^{+}$ denotes the class conditional distribution of positive samples that have the same label as x, and ${p}^{-}$ represents the class conditional distribution of negative samples that have a different label from x. Inspired by Eq (2), we also add $w_{i}$ to Eq (3) to reduce the impact of loss values associated with those AUs that are more commonly activated in the training set.
The contrastive loss in Eq (3) distinguishes itself from the binary cross-entropy loss in Eq (1) in the following aspects: (1) The binary cross-entropy loss is computed based on the sample and its corresponding label y, requiring the encoding of pixel-level information in order to align with the label. In contrast, the contrastive loss does not involve explicit labels but is computed based on the features of samples from different classes. This incentivizes the encoder to capture the differential information among different class samples, without the need to encode pixel-level information, enabling a lightweight encoder. (2) The binary cross-entropy loss constrains the absolute values of the predicted probabilities, while the contrastive loss constrains the difference in similarity between positive and negative sample pairs, without specifying the exact value. This is because the minimal value of $\mathcal{L}_{wnce}$ is 0, which is achieved when the difference between $\text{sim}((f(x),f(x^{+}))$ and $\text{sim}(f(x),f(x^{-}_{j}))$ approaches positive infinity for all $j\in{1,2,\cdots,N}$.
The contrastive learning encoder in our paper is shown in Fig. 2. Different image pairs are input into our backbone, which adopts Swin-transformer [50]. To ensure the fairness of the experimental comparison, The AUs Relationship-aware Node Feature Learning (ANFL) module the same as the encoder for the first stage training of baseline MEGraph [51]. Therefore, because of our discriminative contrastive learning paradigm, our encoder not only has half the parameters less than MEGraph, but also has better experimental results.
III-B Importance re-weighting strategy
To address the second challenge of the class imbalanced problem, we employ a importance re-weighting strategy to minority AUs. This helps to balance the representation of different AUs in the training data and improves the ability of models to accurately classify samples from all AUs. The improved loss ${L}_{aunce}$ is formulated as:
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))^{\beta}}]\}$
@@ III-B Importance re-weighting strategy
where $\beta\in(0,+\infty)$ is the importance weight. Without loss of generality, we consider the similarity calculation as $\text{sim}(f(x),f(x^{+}))=\exp(f(x)^{T}f(x^{+})/\tau)$, where $\tau$ is the temperature coefficient. The inclusion of importance weights transforms the gradient of similarity scores of positive pairs $\text{sim}(f(x),f(x^{+}))$ w.r.t $\mathcal{L}$ into the following form:
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{+}))}=-\frac{1}{\tau}\sum_{j=1}^{N}P_{j}$ (5)
@@
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{-}))}=\frac{1}{\tau}P_{j}$ (6)
So, the relative gradient magnitudes of positive pairs and negative pairs can be expressed as follows:
$|\frac{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{+}))}{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{-}))}|=\frac{\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}$ (7)
When $x$ belongs to the majority class, its positive sample is from the same majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by minority class samples. On the other hand, when x belongs to the minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter update process, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. Theoretically, as long as $\beta_{minority}>\beta_{majority}$, the impact of the imbalanced problem can be mitigated.
III-C Enhanced Supervision Signal
Regarding the third challenge of AU labels tainted by noise and errors, we employ a combined approach that integrates label noisy learning and self-supervised learning techniques to obtain enhanced supervised signals for positive samples. This approach mitigates the adverse effects of noisy labels by introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the model’s robustness to variations and errors present in the labeled dataset.
@@ III-C1 Supervised Signal
Our supervised signal comes from label noisy learning, which aims to address the issue of mislabeled or incorrectly annotated data by introducing mechanisms that account for label noise during training. One common technique is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. Label smoothing allows the model to learn from the noisy labels while reducing the negative impact of label errors. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples. This memory effect is independent of the choice of neural network architecture. Inspired by this observation, for a given sample $x$ and M positive samples $\{x_{i}\}_{i=1}^{M}$ with noisy labels that belong to the same class, we compute the similarity scores between the representation of sample $x$ and the representations of all M positive samples. Based on these similarity scores, we select the positive feature representation with the highest similarity to sample $x$:
$f(x^{+})=f(x_{i}),\text{where}~{}i=\arg\max_{i}f(x)^{T}f(x_{i})$ (8)
@@ III-C2 Self-supervised Signal
$f(x^{+})=f(\mathcal{T}(x))$ (9)
where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{1}^{+}$ and $x_{2}^{+}$, which we consider as a positive pair. Follow SimCLR [21], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to the original size, random color distortions, and random Gaussian blur.
Our third method to obtain positive examples involving a mixture of positive feature representations, which can be seen as a mixup of positive samples to obtain the most representative positive instances.
$f(x^{+})=\frac{1}{M}\sum_{i=1}^{M}f(x_{i})$ (10)
The positive feature representation $f(x^{+})$ obtained from the equation above represents the class centroid of M positive examples $\{x_{i}\}_{i=1}^{M}$. In other words, it synthesizes a virtual representative positive sample, which is also referred to as feature augmentation. This method helps overcome the challenge of instance-level contrast, where each instance is treated as a separate class. By synthesizing a virtual sample that represents the class centroid, the feature representation of each instance is encouraged to be closer to the class centroid, which is benefit of capturing the essential characteristics of the positive class.
The aforementioned three methods of obtaining positive examples combine the use of label noisy learning for supervised signals and data augmentation for self-supervised signals. These three types of positive examples are randomly mixed with probabilities $p_{1}$, $p_{2}$, and $p_{3}$ respectively, where $p_{1}+p_{2}+p_{3}=1$. Our approach aims to improve the quality of supervised signals utilized in the training process. By integrating these techniques, we effectively mitigate the impact of AU labels tainted by noise and errors, increase the resilience of models to errors and variations in the labeled datasets, and ultimately enhance the generalization ability of models to unseen data. The pseudocode for the algorithm is provided below:
@@
By addressing these three challenges, our proposed approach aims to improve the encoding capability of the shared encoder, enhance classification accuracy by addressing class imbalanced problem of each AU type, and enhance the robustness of models to AU labels tainted by noise and errors, ultimately leading to improved generalization performance in AU detection.
@@ IV-A1 Datasets
In our experiments, we extensively utilize two well-established datasets for AU detection, namely BP4D [52] and DISFA [53]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations. For both datasets, we employ a three-fold cross-validation methodology for rigorous evaluation.
BP4D [52]: This dataset comprises 41 participants, including 23 females and 18 males. Each individual is featured in 8 sessions, recorded using both 2D and 3D videos. The dataset encompasses approximately 140,000 frames, each annotated with AU labels indicating occurrence or non-occurrence. Additionally, each frame is meticulously annotated with 68 landmarks detected using the SDM [54] method.
DISFA [53]: This dataset is composed of 27 videos, recorded from 15 men and 12 women. Each video comprises 4,845 frames, resulting in a total of 130,788 images. Frames in this dataset are annotated with AU intensities ranging from 0 to 5, and 68 landmarks are detected using the AAM [55] method. Frames with intensities equal to or greater than 2 are considered positive, while others are classified as negative.
@@ IV-A2 Implementation Details
For the training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of $10^{-5}$ and 60 epochs. Due to our discriminative contrastive learning framework that sample pairs are selected in a same batch, so the batch size has a strong influence on the experimental results. The maximum batch size we can set is 32. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 8 for DISFA and 12 for BP4D respectively, the number K for choosing nearest neighbors in the ANFL is set to 4, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. V-C2.
In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [22, 56, 57]. This involves training a linear classifier atop the fixed base network, and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.

### v3

@@ III-A Contrastive learning framework
III-B Negative sample re-weighting strategy
To address the second challenge of the class imbalance issue, we employ a negative sample re-weighting strategy to minority AUs. This also contributes to hard negative sample mining. Inspired by HCL [58], we apply the following weighting scheme to each negative sample $x_{j}^{-}$:
$\omega_{j}=\text{sim}(f(x),f(x^{-}_{j}))^{\beta},$ (3)
where $\beta\in(0,+\infty)$ can be viewed as the importance weight if we assume that the negative samples follow a von Mises-Fisher (vMF) distribution. The resulting AUNCE loss is:
$\mathcal{L}_{aunce}=\mathbb{E}_{\begin{subarray}{c}{x}\sim{p}_{d}\\ {x}^{+}\sim{p}^{+}\\ {x}^{-}\sim{p}^{-}\end{subarray}}\{-\frac{1}{n_{au}}\sum_{i=1}^{n_{au}}w_{i}$ (4)
$[-\log\frac{\text{sim}(f(x),f(x^{+}))}{\text{sim}(f(x),f(x^{+}))+\sum_{j=1}^{N}\text{sim}(f(x),f(x^{-}_{j}))^{\beta+1}}]\}.$
@@ III-B Negative sample re-weighting strategy
Next, we analyze how this weighting scheme addresses the class imbalance issue and facilitates the mining of hard negative samples. Without loss of generality, we consider the similarity calculation as $\text{sim}(f(x),f(x^{+}))=\exp(f(x)^{T}f(x^{+})/\tau)$, where $\tau$ is the temperature coefficient. The inclusion of importance weights transforms the gradient of $\mathcal{L}$ w.r.t similarity scores of positive pairs into the following form:
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{+}))}=-\frac{1}{\tau}\sum_{j=1}^{N}P_{j},$ (5)
@@
$\frac{\partial\mathcal{L}}{\partial\text{sim}(f(x),f(x^{-}))}=\frac{1}{\tau}P_{j}.$ (6)
Therefore, the ratio between the gradient magnitude of negative examples and positive examples is given by:
$|\frac{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{-}))}{\partial\mathcal{L}/\partial\text{sim}(f(x),f(x^{+}))}|=\frac{\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}{\sum_{j=1}^{N}\exp(f(x)^{T}f(x^{-}_{j})/\tau)^{\beta}}.$ (7)
Above equation follows the Boltzmann distribution. As the value of $\beta$ increases, the entropy of the distribution decreases significantly [59], leading to a sharper distribution in the region of high similarity. This has two consequences: Firstly, it imposes larger penalties on hard negative samples that are in close proximity to the anchor data point, enabling the model to learn more precise decision boundaries from subtle feature differences when AUs activate. Secondly, it controls the relative contribution of negative samples to the parameter updates, providing a means to address class imbalance issues.
Specifically, when anchor data point $x$ belongs to the majority class, its positive sample belongs to the majority class, while its negative sample belongs to the minority class. In this case, we aim to have smaller gradient magnitudes for the positive samples (majority class) to update fewer parameters, while the negative samples (minority class) should have larger gradient magnitudes to update more parameters. To achieve this, we begin to set $\beta\geq 1$, which increases the impact of parameter updates caused by negative samples (minority class). On the other hand, when $x$ belongs to minority class, we begin to set $\beta\leq 1$ to reduce the influence of parameter updates caused by majority class samples. By adjusting the value of $\beta$, we can control the relative importance of positive and negative samples during the parameter updating, ensuring that the model pays more attention to minority class samples when necessary and balances the gradient magnitudes accordingly. This helps to balance the representation of different AUs in training set and improves the model ability to precisely classify all AU samples.
Figure 3: Illustration of our proposed positive sample samplying strategy.
III-C Positive sample samplying strategy
Regarding the third challenge of AU labels tainted by noise and errors, we employ a combined approach that integrates label noisy learning and self-supervised learning techniques to obtain enhanced supervised signals for positive samples. As shown in Fig. 3, this approach mitigates the adverse effects of noisy labels by simultaneously introducing a controlled level of self-supervised signal and supervised signal, thereby enhancing the robustness of model to variations and errors present in the labeled dataset.
@@ III-C1 Supervised Signal
Our supervised signal comes from label noisy learning, which aims to address the issue of mislabeled data by introducing mechanisms that account for label noise during training. One common technique is label smoothing, where a small amount of uncertainty is incorporated into the training process by replacing one-hot labels with smoothed probability distributions. Label smoothing allows the model to learn from the noisy labels while reducing the negative impact of label errors. However, it is challenging to directly apply label smoothing to contrastive learning as the computation of contrastive loss does not involve explicit labels. To overcome this challenge, we turn to the memory effect of neural networks. Previous research has shown that neural networks tend to first fit clean samples and then memorize the noise in incorrectly labeled samples [60, 61, 62]. This memory effect is independent of the choice of neural network architecture. The positive samples that are easier fit are those that are closer to the anchor point and have higher similarity. Inspired by this observation, for a given sample $x$ and M positive samples ${x}_{i=1}^{M}$ with noisy labels belonging to the same class, we calculate the similarity scores between the representation of sample $x$ and the representations of all M positive samples. We then select the positive feature representation with the highest similarity to sample $x$:
$f(x_{h}^{+})=f(x_{i}),\text{where}~{}i=\arg\max_{i}f(x)^{T}f(x_{i}).$ (8)
@@ III-C2 Self-supervised Signal
$f(x_{s}^{+})=f(\mathcal{T}(x)),$ (9)
where $\mathcal{T}(\cdot)$ is a stochastic data augmentation module that applies random transformations to a given data example $x$, generating two correlated views referred to as $x_{s1}^{+}$ and $x_{s2}^{+}$, which we consider as a positive pair. Follow SimCLR [22], we employ a sequential application of three simple augmentations: random cropping and flipping followed by resizing back to the original size, random color distortions, and random Gaussian blur.
Our third method to obtain positive examples involving a mixture of positive feature representations, which can be seen as a mixup of positive samples to obtain the most representative positive instances:
$f(x_{m}^{+})=\frac{1}{M}\sum_{i=1}^{M}f(x_{i}).$ (10)
The positive feature representation $f(x_{m}^{+})$ obtained from the equation above represents the class centroid of M positive examples $\{x\}_{i=1}^{M}$. In other words, it synthesizes a virtual representative positive sample, which is also referred to as feature augmentation. This method helps overcome the challenge of instance-level contrast, where each instance is treated as a separate class. By synthesizing a virtual sample that represents the class centroid, the feature representation of each instance is encouraged to be closer to the class centroid, which is benefit of capturing the essential characteristics of the positive class.
The aforementioned three methods of obtaining positive examples combine the use of label noisy learning for supervised signals and data augmentation for self-supervised signals. These three types of positive examples are randomly mixed with probabilities $p_{1}$, $p_{2}$, and $p_{3}$ respectively, where $p_{1}+p_{2}+p_{3}=1$. Our approach aims to improve the quality of supervised signals utilized in the training process. By integrating these techniques, we effectively mitigate the impact of noisy and false labels, increase the resilience of models to errors and variations in labeled datasets, and ultimately enhance the generalization ability of models to unseen data.
By addressing the above three challenges, our approach improves the encoding capability of the shared encode, ultimately leading to improved generalization performance in AU detection. The pseudocode for the algorithm is provided in Algorithm. 1.
@@ IV-A Experiment Setting
TABLE I: AU occurrence rates (%) in the training sets of different datasets. ”-” denotes the AU is not annotated in this dataset.
Dataset AU Index
1 2 4 6 7 9 10 12 14 15 17 23 24 25 26
BP4D [63] 21.1 17.1 20.3 46.2 54.9 - 59.4 56.2 46.6 16.9 34.4 16.5 15.2 - -
DISFA [64] 5.0 4.0 15.0 8.1 - 4.3 - 13.2 - - - - - 27.8 8.9
GFT [65] 3.7 13.5 3.7 28.3 - - 24.6 29.3 3.1 10.7 - 25.0 14.1 - -
Aff-Wild2 [66] 11.9 5.1 16.0 26.5 39.9 - 34.5 24.3 - 2.8 - 3.1 2.8 62.8 7.6
@@ IV-A1 Datasets
In our experiments, we extensively utilize three well-established datasets in constrained scenarios for AU detection, namely BP4D [63], DISFA [64], GFT [65], and one well-established dataset in unconstrained scenarios, namely Aff-Wild2 [66]. Our primary focus is on assessing frame-level AU detection, as opposed to datasets that solely provide video-level annotations.
BP4D [63]: This dataset comprises 41 participants, including 23 females and 18 males. Each individual is featured in 8 sessions, recorded using both 2D and 3D videos. The dataset encompasses approximately 140,000 frames, each annotated with AU labels indicating occurrence or non-occurrence. Additionally, each frame is meticulously annotated with 68 facial landmarks detected using the SDM [67] method. For BP4D dataset, we employ a three-fold cross-validation methodology for rigorous evaluation.
DISFA [64]: This dataset is composed of 27 videos, recorded from 15 men and 12 women. Each video comprises 4,845 frames, resulting in a total of 130,788 images. Frames in this dataset are annotated with AU intensities ranging from 0 to 5, and 68 facial landmarks are detected using the AAM [68] method. Frames with intensities equal to or greater than 2 are considered positive, while others are classified as negative. We also employ the same three-fold cross-validation methodology as BP4D dataset for rigorous evaluation of DISFA dataset.
GFT [65]: This dataset comprises 96 participants, organized into 32 groups of three individuals, engaged in unscripted conversations. Each participant is recorded via video, with the majority of frames capturing moderate out-of-plane poses. Annotations for each frame include 10 AUs, alongside 68 facial landmarks identified using the ZFace [69] tool. In accordance with the standard training and testing partitions, our training set consists of 78 participants, amounting to approximately 108,000 frames, while the testing set includes 18 participants, totaling around 24,600 frames.
Aff-Wild2 [66]: This dataset is an extensive in-the-wild facial expression dataset, with videos sourced from YouTube, encompassing a broad range of ages, ethnicities, professions, emotions, poses, lighting conditions, and occlusions. This dataset comprises 305 videos, totaling approximately 1,390,000 frames for the training set, and 105 videos, with around 440,000 frames for the validation set. Each frame is annotated with 12 AUs. For facial landmark annotations, we utilize the Dlib library [70], [71] for detecting 68 facial landmarks per frame. Following the methodology of Zhang et al. [72], we train our AUNCE on the training set and evaluate it on the validation set.
The occurrence rates of AUs in the training sets of these datasets are presented in Table I. It is evident from the table that the class imbalance issue of each type of AU is prevalent across all datasets, especially on the DISFA, GFT and Aff-Wild2 datasets.
@@ IV-A2 Implementation Details
For the training process, the encoder is conducted on one NVIDIA GeForce RTX 3090 GPU with 24GB for one day and implemented by PyTorch using an AdamW optimizer with $\beta_{1}$= 0.9, $\beta_{2}$= 0.999 and weight decay of $10^{-6}$. We set a learning rate of $10^{-5}$ and 60 epochs. Due to our discriminative contrastive learning framework that sample pairs are selected in a same batch, so the batch size has a strong influence on the experimental results. In terms of the hyperparameters in AUNCE, $n_{au}$ is set as 12 for BP4D and 8 for DISFA respectively, the number K for choosing nearest neighbors in the ANFL [56] is set to 4, and the temperature coefficient $\tau$ is 0.5. The analysis for $\beta$ and $p_{1}$, $p_{2}$, $p_{3}$ will introduce in Section. IV-D2.
In assessing the acquired representations, we adhere to the established linear evaluation protocol, as commonly employed in previous studies [27, 73, 74, 22]. This involves training a linear classifier atop the fixed base network, and the highest test accuracy across our models serves as a surrogate measure for the quality of the learned representations.